The standard cell potential (E°) of a voltaic cell constructed using the cell reaction below is 0.76 V:

Zn(s) + 2H+(aq) → Zn2+(aq) + H2(g)

With PH2= 1.0 atm and [Zn2+] = 1.0 mol L-1, the cell potential is 0.37 V. The concentration of H+ in the cathode compartment is ________ mol L-1.

Answers:

2.6 × 10-7

5.1 × 10-4

0.686.6 × 10-14

4.3 × 10-27
Answer #1

Zn(s) + 2H+ Zn2+ + H2(g)

Number of electrons transferred (n) = 2

[ Zn] = 1

Then

Nernst equation is

Ecell = Eocell - log{ [Zn2+] P(H2) / [H+] }

0.37 = 0.76 - log (1*1)/[H+]2

Or, (0.37 - 0.76) = log [H+]2

Or, - 0.39 = 0.0295 log [ H+ ]2

or, log [H+]2 = (0.39/0.0296)

Or, 2 log [H+] = - 13.17

Or, log [H +] = - 6.587

Or, [H+] = 10-6.587 = 2.6 *10-7 M

Hence, concentration of H+ = 2.6*10-7 mol/L.